solve each equation for the given variable


a) b - 8 = -4 b) 25x + 2 = 77 c) k + 7 = -4 d) 5w/3 = 15 ( this is 5w over 3)


1
Expert's answer
2021-05-03T04:52:02-0400

(a) "b-8=-4\Rightarrow b=8-4=4"


(b) "25x+2=77\Rightarrow 25x=77-2\Rightarrow 25x=75\Rightarrow x=3"


(c) "k+7=-4\Rightarrow k=-7-4=-11"


(d) "\dfrac{5w}{3}=15\Rightarrow w=\dfrac{15\times 3}{5}=9"